OpenStudy (anonymous):

Highest common factor of 9 and 15????

OpenStudy (anonymous):

3

myininaya (myininaya):

9=3*3 15=3*5

myininaya (myininaya):

they only have a 3 in common

OpenStudy (anonymous):

exactly whatever she said LOL

OpenStudy (anonymous):

omg im so stupid, probably cause im reallllyy tired :D thanks again

myininaya (myininaya):

omg you called me a she

myininaya (myininaya):

usually people just assume i;m a guy

myininaya (myininaya):

you get a medal

OpenStudy (anonymous):

I thought myininaya was a girl name sorry man.

myininaya (myininaya):

you aren't stupid jordan

myininaya (myininaya):

no you were right

myininaya (myininaya):

lol i just couldn't believe it

OpenStudy (anonymous):

sorry i gave you a medal for my mistake lol

OpenStudy (anonymous):

by the way answer my question if you can.

myininaya (myininaya):

no there was no mistake

myininaya (myininaya):

lol

OpenStudy (anonymous):

how can you look at my question just by loooking at someones profile?

myininaya (myininaya):

i think you can just click on the question from someone's profile are you asking how to get to one's question from the profile?

myininaya (myininaya):

i have to go to class but i will look at your question later k? :)

OpenStudy (anonymous):

yeah like for example i want for you to help me with a question right now. but i wanna know how do you do that so you can see my question. u know what i mean?

OpenStudy (anonymous):

in other words what do you click so you can see my question that i just posted up a while ago... or can I just post it here. but then i cant give you any medals and stuff lol

myininaya (myininaya):

your question is right below this one

myininaya (myininaya):

if you become my fan i can look at your profile and look at your question from earlier i mean even if you didn't become my fan i can search through all the questions in hopes of finding yours if someone doesn't answer your question soon you should repost it though because usually people don't go through all of that to find questions to answer

myininaya (myininaya):

i have to go. i'm sorry peace